OpenStudy (anonymous):

if f(x) = 3x^2 - 2 , find the inverse and list any restrictions on the domain and range required to make it the inverse. help asap plz!!!

Parth (parthkohli):

Hint: Make it \(y = 3x^2 - 2 \Longrightarrow x = 3y^2 - 2\) and solve for \(y\).

OpenStudy (anonymous):

hey can u show me the work please ? since im just reviewing for my final exam

OpenStudy (anonymous):

You have: \[\large x = 3y^2 - 2\] Now add 2 both the sides and tell me what did you get..

Parth (parthkohli):

You must do the work yourself, Abbey. I may give hints and examples.

OpenStudy (anonymous):

2x=3y2-4?

Parth (parthkohli):

For example, you may find out \(f^{-1}(x)\) in this case:\[ \color{Black}{\Rightarrow f(x) = 3 + x } \Longrightarrow x = 3 + y \Longrightarrow {x - 3 } = y \Longrightarrow f^{-1}(x) = x - 3\]

Parth (parthkohli):

Nope. What did you do here?

OpenStudy (anonymous):

im lost .

Parth (parthkohli):

Where are you lost?

OpenStudy (anonymous):

nick?

OpenStudy (anonymous):

f(x)=y y=3x^2-2 sqrt(y-2)/3=x y>2

OpenStudy (anonymous):

f^-1(x)=sqrt(y-2)/3

OpenStudy (anonymous):

its a pre cal question